Please help me in my hw :(

A 25-liter solution of alcohol in water is 72% alcohol. How many liters of an 80% solution must be added if the final solution is to be 75% alcohol?
Big thanks :*

Answer by checkley79(3341)

.72*25+.80X=.75(25+X)
18+.80X=18.75+.75X
.80X-.75X=18.75-18
.05X=.75
X=.75/.05
X=15 LITERS OF 80% SOLUTION IS REQUIRED.
PROOF:
.72*25+.80*15=.75(25+15)
18+12=.75*40
30=30